OpenStudy (nesha97):

9. Write a function with the given characteristics. A polynomial with rational coefficients having roots 3, 3, and 3 − i Help Please

OpenStudy (eliesaab):

If it has 3-i as a root, then it will have 3+i as a root

OpenStudy (eliesaab):

Just expand \[ (x-3)^2 (x-(3+i)) (x-(3-i)) \]

OpenStudy (eliesaab):

Are you here

OpenStudy (eliesaab):

For example \[ (x-3)^2=x^2-6 x+9 \] That should be easy

OpenStudy (eliesaab):

Also notice that \[ (x-(a-i b)) (x-(a+i b))=((x-a)-i b) ((x-a)+i b)=\\(x-a)^2-(i b)^2=a^2-2 a x+b^2+x^2 \]

OpenStudy (eliesaab):

Use the above to expand \[ (x-(3+i)) (x-(3-i) \]

OpenStudy (eliesaab):

Then multiply the two expansions above and you will get your polynomial

OpenStudy (nesha97):

sorry and im confused